### v1

# The Unseen Species Problem Revisited

Abstract
The unseen species problem is a classical problem in statistics. It asks us to, given $n$ i.i.d. samples from an unknown discrete distribution over an unknown set, predict how many never before seen outcomes would be observed if $m$ additional samples were collected. For small m we show the classical but poorly understood Good-Toulmin estimator to be minimax optimal to within a factor $2$ and resolve the open problem of constructing principled prediction intervals for it. For intermediate m we propose a new estimator which achieves the minimax error for linear estimators up to an explicit multiplicative constant. Our estimator vastly outperforms the standard Smoothed Good-Toulmin estimator in the worst case and performs substantially better on several real data sets, namely those with many rare species. For large m we show that a previously mentioned estimator which did not have known rate guarantees actually achieves a marginally better rate than subsequent work. We find that this marginal rate improvement translates to meaningfully better performance in practice. We show in all three regimes that the same methods also achieve the same rate on incidence data, without further independence assumptions, provided that the sets are of bounded size. We establish, by means of bounded size biased couplings, concentration for some natural functionals of sequences of i.i.d. discrete-set-valued random variables which may be of independent interest.
1 Introduction
In the early 1940s the British naturalist Alexander Corbet spent two years trapping butterflies in British Malaya [20]. Upon his return to England he consulted with R.A. Fisher as to how many new species he may expect to discover on a return expedition [20]. Making such a prediction has become known as the unseen species problem. With access to a good estimate one could, for example, make a more informed decision about whether a return trip is worthwhile.
The return expedition needs not be of the same length as the first expedition. In the limit of an infinitely long return expedition the unseen species problem becomes the classical support size estimation problem. At the other extreme, namely a return expedition consisting of capturing only a single specimen, it becomes the classical missing mass problem in which one tries to estimate how much probability mass is associated to as of yet unobserved outcomes. In this way the unseen species problem ”interpolates” between these classical problems, each arising as an extremal version of the unseen species problem. As we feel the two are sometimes confused, we spend a few more sentences on distinguishing the unseen species problem from support size estimation. Support size estimation is asking ”how much could one discover?” and the unseen species problem is asking ”how much will we discover?”. For example, in the context of software code [6] support size estimation is estimating the number of bugs that exist, while the unseen species problem is the question of determining how many bugs will be encountered in some finite time interval. As another example, there were a large number of body fragments collected in the aftermath of the 9/11 attacks, so many in fact that as of the writing of [1] the identification effort was still ongoing. If one imagines that each fragment is sampled i.i.d. from some distribution over the victims, then the support size is the number of victims, which is known by other means [1]. The task carried out in the paper, namely to predict the number of victims who are represented in the finite number of recovered fragments, is the unseen species problem.
1.1 The Generalized Unseen Species Problem
Let $\mathcal{S}$ be the unknown, but presumed at most countable, set of species. Following [15] we let $n,m$ be Poisson random variables with means $t,t^{\prime}$, respectively. At least in the classical problem $t$ and $t^{\prime}$ are naturally interpreted as the duration of the expeditions, with $n$ and $m$ being the number of specimen caught. In the classical unseen species problem one takes the observations $X_{1},\ldots,X_{n}$ to be i.i.d. samples from an unknown $\mu\in\mathcal{P}(\mathcal{S})$, where $\mathcal{P}(\cdot)$ denotes the set of probability measures over a set, and attempts to predict how many new outcomes will be observed in the next $m$ samples. We consider the following generalization: Let the data consist of $X_{1},\ldots,X_{n}$, i.i.d. samples from an unknown $\mu\in\mathcal{P}\left(\mathcal{T}(\mathcal{S})\right)$, where $\mathcal{T}(\mathcal{S})$ denotes the set of finite multisets over $\mathcal{S}$. For each species $s$, let $N_{s}$ be the number of times it was observed, counted with multiplicity. Let $\mathcal{S}_{t}:=\{s\in\mathcal{S}:N_{s}>0\}$ denote the set of species observed until time $t$, $S_{t}:=|\mathcal{S}_{t}|$ and define $T:=t+t^{\prime}$. We then call the task of predicting
$S_{t,T}:=\left|\mathcal{S}_{T}\setminus\mathcal{S}_{t}\right|=S_{T}-S_{t},$
the unseen species problem.
The classical unseen species problem is recovered by insisting that $\mu$ is supported on singletons. We hope that our model better captures the dependence structure in the problem, which we give some examples to illustrate.
Example 1.
When collecting butterflies one expects clustering of members of the same species and of species which enjoy the same micro-environments. It is therefore much more plausible to believe that every day a set of butterflies is caught and that independence arises between days rather than individuals.
Example 2.
To sequence a genome from a single cell one prepares a ”library” of DNA fragments which haphazardly cover different parts of the target genome. Then a machine reads those fragments. One would like to predict from a small number of fragments how much of the genome could be recovered from reading many fragments from that library as then one could compare protocols for preparing libraries without having to do (expensive) deep sequencing experiments [14]. Rather than genomic locations being independent, it is much more plausible that fragments are independent.
Example 3.
Suppose that an intelligence agency is spying on a terrorist organization by means of monitoring phone calls and is facing a decision: either arrest the known terrorists at risk of sending the rest into hiding, or wait to discover more terrorists. Clearly being able to forecast the number of new terrorists that will be discovered by waiting is key to making an informed decision. The usual model would have one believe that who is one end of a phone call is independent of who is at the other end. It is much more plausible to believe that calls rather than callers are independent.
Example 4.
In many trading card games customers can buy small packages (”packs”) of randomized cards. The customer may wish to forecast the growth of their collection as they buy such packs, which is closely related to the famous coupon collection problem. The number of cards in a pack is known and fixed but the exact contents are random and unknown. The cards are not independent as various constraints are imposed. For example, the presence of precisely one rare card, roughly equal presence of cards of various types etc. Hence it is much more plausible to believe that packs rather than cards are independent. We thank Torben Swoboda who thought of this application.
That data is often structured in ways resembling the above examples is well-known to biologists who refer to it as incidence data [9] and to statisticians who speak of feature allocation models [2] or edge exchangeable random graphs [13]. Edge exchangeable random graphs are a way to encode the data as a multi-hyper-graph by setting the observed species as the vertices and the observed sets as the edges. However, for estimation problems resembling those discussed here the only analyses we are aware of assume that membership in each set is independent for each species [9][2], essentially assuming away precisely the dependence we are trying to capture. We will often assume that the sets are of bounded size, but make no such within-set independence assumption. Owing to the De-Finetti type result of [13] one could assume that the data is an exchangeable sequence of sets rather than an independent one with suitable modifications to our results. Concretely, the De-Finettti type result says, essentially, that weakening independence corresponds to allowing $\mu$ itself to be random. However, many of our bounds are worst case bounds which automatically hold also for random $\mu$.
1.1.1 Reduction to Multiplicity-Free Case
When we observe multisets we may forget about the duplicates without changing when we discover what species, in particular, without changing $S_{n,n+m}$. Thus in the remainder it suffices to consider $\mu\in\mathcal{P}(2^{\mathcal{S}})$ (still supported on finite sets). Perhaps it appears that we have robbed ourselves of some information that our estimators [^1] Following previous authors we use the term ”estimator” even though ”predictor” is perhaps more accurate. could depend on, but note that starting with a $\mu\in\mathcal{P}(2^{\mathcal{S}})$ we can add multiplicities arbitrarily without changing when what is discovered, so in general the duplicates cannot be informative.
Remark 5.
As observed in [8], while such a reduction is always possible it is sometimes near mandated. Therein a data set of tropical fish is considered. The data is collected by groups of divers submerging and then, upon their return to the surface, reporting what they observed. As noted in [8] there is a substantial risk that the same specimen was seen by multiple divers causing essentially ”illusory” duplicate observations. Additionally, counting the number of fish in a school may be hopeless.
1.2 Previous Work and Our Contribution
Different methodologies are suitable for different ranges of $r:=\frac{t^{\prime}}{t}$. To reflect this we speak of the near, intermediate and distant future, corresponding roughly to $r\leq 1$, $1<r\lesssim\log(t)$ and $\log(t)\lesssim r\lesssim\exp(t^{\frac{\alpha}{2}})$ for a constant $\alpha\in(0,1)$, respectively. Before describing the literature and our contribution more accurately, we give a crude overview in Table 2.
Near Intermediate Distant
Classical EB $\star$ RO $\star$ AG $\star$ EB $\checkmark^{\prime}$ RO $\checkmark^{\prime}$ AG $\star$ EB $\star$ RO $\checkmark^{\prime}$ AG $\star$
Generalized EB $\star$ RO $\star$ AG $\times$ EB $\star$ RO $\star$ AG $\times$ EB $\star$ RO $\star$ AG $\times$
Table 1: Overview of relevant results, categorized in three prediction ranges (near, intermediate, distant) and two set-ups (classical, generalized). EB: Non-asymptotic error bounds, RO: Rate and optimality, AG: Asymptotic Gaussianity. $\checkmark^{\prime}$: We sharpen a pre-existing result. $\star$: Our contribution. $\times$: Still an open problem [^2] Even showing that $S_{t}$ is asymptotically Gaussian, never mind the error in predicting it, is an open problem, even when the sets are of size two. See problem 6.9 of [18] and progress in [16].
1.2.1 Near Future
For near future prediction the classical Good-Toulmin estimator (GTE),
$\hat{S}_{t,T}^{(GT)}:=-\sum_{s\in\mathcal{S}}1_{N_{s}>0}(-r)^{N_{s}},$
has long been the dominant approach but little was known about it theoretically. We show that for any $r\in[0,1]$ and $t>0$ the worst case average error of the GTE is within a factor of two from the worst case average error achievable by any eligible estimator. We show that no other linear estimator can achieve anything but the most trivial of performance gains on the GTE. We show that no other estimator which respects some natural problem symmetries other than precisely the GTE achieves a rate which is better than that of the estimator which always predicts zero, henceforth referred to as the null-estimator. We define a variance proxy and show that a central limit theorem holds with this variance proxy, settling the open problem [11] of constructing principled prediction intervals. Practitioners [1] have previously estimated $\mu$ and then simulated the unseen species problem on $\hat{\mu}$ [9] which lacks theoretical guarantees.
In the generalized problem we explicitly compute the extra error incurred by moving from the classical to the generalized setting. We show generalizations of all the minimax results to the generalized unseen species problem with sets of size bounded by $B$. We show that the extra error can be upper bounded not just proportionally to $B^{2}t$ (which gives the optimal rate) but in terms of the typically much smaller $B\mathbb{E}[S_{t}]$. This means we get the expected rate not just in a uniform sense but also in $\mathbb{E}[S_{t}]$, which is a much more delicate result.
1.2.2 Intermediate Future
In the intermediate future the variance of the GTE is so large as to make it unusable and one must seek other estimators. The most well-known alternative are the Smoothed Good-Toulmin estimators (SGTE) proposed by [20], which encompass the Efron-Thristed estimator of [15]. They are linear estimators which work by averaging over stochastic truncations of the GTE. Specifically, if $L$ non-negative-integer-valued random variable then
$\hat{S}_{t,T}^{(L)}=-\sum_{s\in\mathcal{S}}1_{N_{s}>0}P(L\geq N_{s})(-r)^{N_{s}},$
is the SGTE (SGTE) associated to (the distribution of) $L$. In a suitably normalized metric such estimators achieve the rate $n^{-\frac{1}{r}}$ [20]. There exists a constant $c$ so that in the same metric one cannot obtain a rate better than $n^{-\frac{c}{r}}$ [20]. This means that SGTEs are minimax optimal ”up to constants” [20], but because there is a constant in the exponent, their result still leaves room for meaningful improvement and we speak of their rates being pseudo-optimal.
We give a recipe, in terms of a certain optimization problem, for constructing linear estimators which are, up to an explicit multiplicative constant, optimal in the class of linear estimators (and consequently (at least) pseudo-optimal in general owing to the SGT being linear). Algorithmically searching for good estimators hearkens back to [15], but therein they were optimizing a-posteriori while we optimize a-priori. Unfortunately and somewhat mysteriously the proof of the optimality result does not give the rate. Our estimator is much better in the worst case and substantially better on several real data sets, especially when there are many rare species. We give a central limit theorem with a variance proxy for linear estimators but since the estimators are in general biased this does not immediately allow the construction of prediction intervals.
In the generalized problem with sets of bounded size we show quite generically that linear estimators do not suffer a rate degeneration compared to the classical problem. Consequently, the estimators of [20] are psuedo-optimal also in this subset of the generalized problem, with the same rate gap as in the classical problem.
1.2.3 Distant Future
To make predictions beyond the impossibility results of [20] some extra assumption is needed, perhaps the most natural one being a regular variation (power law tail) assumption. Namely, one assumes that for some constants $c>0$ and $\alpha\in(0,1)$,
$\nu(x):=\sum_{s\in\mathcal{S}}1_{\mu(\{s\})>x}\approx cx^{-\alpha}.$ (1)
The consequences of making such an assumption in the unseen species problem are briefly mentioned in [4] and subsequently explored in detail in [17]. From these works it is known that if $\mu$ satisfies (1) and one finds an accurate estimator $\hat{\alpha}$ for $\alpha$ then
$\hat{S}_{t,T}:=S_{t}((1+r)^{\hat{\alpha}}-1),$ (2)
is an accurate estimator in the unseen species problem. The prediction problem is thus reduced to a parameter estimation problem. The authors of [17] gave an estimator for $\alpha$, motivated by certain maximum likelihood considerations, which allowed the induced estimator for the generalized unseen species problem to achieve a rate of $\frac{\log(t)\log(r)^{2}}{t^{\alpha}}$ in a suitably normalized metric. They showed this rate to be optimal up to polylog factors. With $\phi_{1}$ being the number of species seen precisely once, we suggest using $\hat{\alpha}:=\frac{\phi_{1}}{S_{t}}$ instead, which was mentioned and shown to give rise to a multiplicatively consistent unseen species estimator in [4], which predates [17]. We give a pre-asymptotic and explicit tail bound in terms of estimable quantities, allowing the construction of conservative prediction intervals. From it one also sees that the induced unseen species estimator does not suffer the $\log(t)$ factor in the rate. The rate improvement means that we improve the prediction horizon to $r(t)$ growing like $\exp\left(\sqrt{t^{\alpha}}\right)$ rather than $\exp\left(\sqrt{\frac{t^{\alpha}}{\log(t)}}\right)$. Experimentally we find that this somewhat marginal rate improvement translates to substantial improvements on real data. Moreover, despite it being non-linear, its simplicity allows us to show that it achieves the same rate on the generalized problem when the sets are of bounded size, implying that it achieves the minimax rate up to poly-log factors. The proof is via a new concentration inequality for, among other quantities, $S_{t}$ and $\phi_{1}$, for incidence data. It is obtained as an application of the work of [3] on Stein’s method for concentration via sized biased couplings for configuration models. Equivalently it gives concentration for the number of vertices in an edge exchangeable graph and for the number of vertices of a given degree. Such degrees were of interest in [18]. On the Bayesian side it is known that if we one restricts oneself to $\mu$ satisfying (1) then the Pitman-Yor process serves as a reasonable prior and that Gaussian credible intervals can be constructed for this approach [11].
1.3 Structure
The following Section 2 discusses some symmetries it would be desirable for the estimators to display. Then the next three sections, 3, 4, 5 deal with the near future, intermediate future and distant future, respectively. These three sections are each split into two sub-sections, the first covers the classical problem and the second the generalized problem. Finally we compare the estimators empirically in Section 6.
Full proofs are relegated to the appendix but when a proof permits an elucidating, short, high-level summary it is provided as a proof sketch.
2 Linear Estimators and Problem Symmetries
For a species $s\in\mathcal{S}$, recall that $N_{s}$ is the number of observations of $s$ in the sample and let $N_{s}^{\prime}$ be the number of observations of $s$ we will make on the return expedition. For $i\in\mathbb{Z}_{>0}$, let $\phi_{i}:=\sum_{s\in\mathcal{S}}1_{N_{s}=i}$ be the number of species observed exactly $i$ times and encode the sequence as a vector $\boldsymbol{\phi}:=(\phi_{1},\phi_{2},\ldots)$.
Definition 6.
We say that an estimator is atemporal if the estimator is a function only of $r$ and $\boldsymbol{\phi}$, so in particular does not (explicitly) depend on time.
Why atemporality is a natural property is most clear in the classical un-poissonized setting. There $n$, which is analogous to $t$, satisfies $n=\sum_{i=1}^{\infty}i\phi_{i}$. Thus allowing dependence on $n$ adds no extra expressive power.
Definition 7.
We say that an estimator is union-additive if
[…]
for all $\boldsymbol{\phi},\tilde{\boldsymbol{\phi}},t,\tilde{t}$.
To motivate union-additvity, imagine that there are two biologists in locations with disjoint sets of species collecting samples and making predictions for how many new species they will discover. Imagine also that there is a research group leader with access to the union of the data sets. Then the number of discoveries made by the group is the sum of the discoveries made by the researchers. Union-additivity asks that the predictions satisfy the same equality.
Definition 8.
We say that an estimator is $\boldsymbol{\phi}$-homogeneous if $\hat{S}_{t,T}(r,k\boldsymbol{\phi})=k\hat{S}_{t,T}(r,\boldsymbol{\phi}).$
To motivate $\phi$-homogeneity, imagine that we are sampling matching pairs of shoes. We can either see this as the classical unseen species problem with an individual being a pair of shoes or as the generalized unseen species problem with shoes as individuals and all sets of size two (one left and one right shoe, which we consider distinct). Then $\phi$-homogeneity asks that the prediction should not depend on which representation of the problem we choose.
Definition 9.
We say that an estimator is linear if there exists a function $H:\mathbb{Z}_{\geq 0}\mapsto\mathbb{R}$, in general depending on $r$ and $t$ such that,
$\hat{S}_{t,T}=\hat{S}^{(H)}_{t,T}:=\sum_{s\in\mathcal{S}}H(N_{s}),$
and $H(0)=0$.
In our proofs we will also often tacitly assume that $H$ is such that linearity of expectation holds which is not restrictive. The condition $H(0)=0$ is enforced because unobserved species cannot contribute. Notice that
$\sum_{s\in\mathcal{S}}H(N_{s})=\sum_{i=1}^{\infty}H(i)\phi_{i},$
so we could equally well have considered estimators which are linear in the $\phi_{i}$:s, justifying the name. We will often view $H$ as a sequence and so often write $H_{i}$ in place of $H(i)$. The main appeal of linear estimators is that they constitute a large enough class to contain performant estimators, yet is simple enough to reason about. However, the following proposition and remark say that they play well with the symmetries of the problem.
Proposition 10.
If an estimator is atemporal then it is linear if and only if it is union-additive. In particular, if an estimator is atemporal and union-addititve then it is linear.
Remark 11.
Linear estimators are automatically $\phi$-homogeneous. However, $\phi$-homogeneous estimators need not be linear and our suggested estimator in the distant future regime is of this type.
3 Near Future
3.1 Classical Setting
3.1.1 Minimax Optimality of the Good Toulmin Estimator
Our feeling is that the following should be known but the literature seems not to contain the precise statement.
Proposition 12.
For $t\geq 0,r\leq 1$,
$\mathbb{E}[|S_{t,T}-\hat{S}^{(GT)}_{t,T}|^{2}]$ $=\mathbb{E}[S_{t,T}]+\mathbb{E}\left[\sum_{s\in\mathcal{S}}1_{N_{s}>0}r^{2Ns}\right],$
$\sup_{\mu}\mathbb{E}[|S_{t,T}-\hat{S}^{(GT)}_{t,T}|^{2}]$ $=r(r+1)t.$
Let $\mathcal{F}_{t}$ be the filtration generated by the Poisson processes.
Theorem 13.
For all $r\geq 0$ and $t\geq 0$ and all $\mathcal{F}_{t}$-measurable estimators $\hat{S}_{t,T}^{*}$ we have
$\sup_{\mu}\mathbb{E}[|S_{t,T}-\hat{S}_{t,T}^{*}|^{2}]\geq rt.$
Proof sketch.
Computing the error in using the conditional expectation as an estimator. ∎
From this one sees that for $r\leq 1,$ the GT-estimator is minimax optimal to within a factor $1+r\leq 2$. Within the class of linear estimators one can find an even tighter lower bound, ruling out any non-trivial improvement on the GTE within this class.
Theorem 14.
For any linear estimator $\hat{S}^{(H)}_{t,T}$ and any $t,r\geq 0$ we have
$\sup_{\mu}\mathbb{E}[|S_{t,T}-\hat{S}^{(H)}_{t,T}|^{2}]\geq r^{2}\left(\frac{t}{t+1}\right)^{2}+rt+\left(\frac{t}{t+1}\right)^{2}r^{2}t.$
Proof sketch.
We consider the error incurred on a distribution with a very large number of very rare species. This error is determined by $H_{1}$, which we then optimize over. ∎
If we restrict the class even more then we can show a substantial rate gap to any other estimator. We write $f(t)\lesssim g(t)$ to mean $f(t)\leq cg(t)$ for some constant $c$ independent of $t$ and all $t$ large enough.
Theorem 15.
Fix $r>0$ and let $\hat{S}_{t,T}^{(H)}$ be an atemporal and union-additive estimator other than precisely the GTE, then
$\sup_{\mu}\mathbb{E}[|S_{t,T}-\hat{S}^{(H)}_{t,T}|^{2}]\gtrsim t^{2}.$
Proof sketch.
The hypothesis on the estimator allows the bias to be written as a sum over the species with each term depending only on the product $p_{s}t$. Now double $t$ and transform $\mu$ by, for every species creating a new species and transferring half the probability mass from the old species to the new species. For all the old species $p_{s}t$ is unchanged since $p_{s}$ was divided by two and $t$ was multiplied by two. However, the sum now has had each term duplicated and hence is twice as large. From this we see that the bias can be made to grow linearly in time and hence the MSE grows quadratically. The GTE escapes by being unbiased. ∎
Remark 16.
The null-estimator achieves the same rate so either one achieves the optimal rate or the same rate achieved by the null-estimator. Since the GTE is dreadful for $r>1$ this also means that when $r>1$ there is no reasonable estimator which respects the symmetries of the problem. That is to say, in this sense prediction is impossible beyond $r=1$.
3.1.2 Non-uniform Error Theory
The following theorem allows the construction of principled prediction intervals. We estimate the variance by
$\hat{V}_{t}:=\sum_{s}1_{N_{s>0}}r^{2N_{s}}+\hat{S}_{t,T},$
and let $\mathcal{N}(0,1)$ denote the standard normal distribution.
Theorem 17.
If $r\leq 1$ and $\mathbb{V}[S_{t,T}-\hat{S}^{(GT)}_{t,T}$$\rightarrow\infty$ as $t\to\infty$ then
$\frac{S_{t,T}-\hat{S}_{t,T}^{(GT)}}{\sqrt{\hat{V}_{t}}}\rightarrow\mathcal{N}(0,1),$
in distribution as $t\to\infty$.
If $\mathcal{S}$ is finite then there is not really an asymptotic regime to speak of and we have the following.
Proposition 18.
Let $\mu$ be supported on a finite set of species and fix $r<1$. Then almost surely
$S_{t,T}$ $\rightarrow 0,\qquad\qquad\qquad$ $\hat{S}_{t,T}$ $\rightarrow 0,\qquad\qquad\qquad$ $\hat{V}_{t}$ $\rightarrow 0.$
3.2 Generalized Setting
Because the proof relies only on linearity of expectation the GTE remains unbiased in the generalized setting. It transpires that the error in applying it can be seen as the ”usual” (classical) error $\delta(t):=\mathbb{E}[S_{t,T}]+\mathbb{E}[\sum_{s\in\mathcal{S}}1_{N_{s}>0}r^{2N_{s}}]$ and an extra error term arising from the dependence between different species. We define some notation that will be useful for expressing this extra error and other quantities. For $x,y\in\mathcal{S}$ let
$M_{x\cap y}$ $:=\sum_{C\in 2^{\mathcal{S}}:C\supset\{x,y\}}\mu(C)$ $M_{x\cup y}:=\sum_{C\in 2^{\mathcal{S}}:C\cap\{x,y\}\neq\emptyset}\mu(C)$
$M_{x}$ $:=\sum_{C\in 2^{\mathcal{S}}:x\in C}\mu(C),$ $M_{x\backslash y}:=\sum_{C\in 2^{\mathcal{S}}:x\in C,y\notin C}\mu(C).$
We adopt the convention that $\sum_{x,y}$ means the sum over ordered pairs of distinct elements of $\mathcal{S}$. Let
$\epsilon(t):=\sum_{x,y}e^{-(1+r)M_{x\cup y}t}(e^{r(r+1)M_{x\cap y}t}-1).$
Theorem 19.
In the generalized unseen species problem,
$\mathbb{E}[|S_{t,T}-\hat{S}^{(GT)}_{t,T}|^{2}]=\delta(t)+\epsilon(t).$
It may be that $\epsilon(t)$ is not small. To demonstrate this we give an example showing that the generalized unseen species problem is not always feasible.
Example 20.
Suppose all animals live in a single cave such that either none or all of them are discovered. This corresponds to a $\mu$ which assigns positive probability only to a single large set. By making this set arbitrarily large and tuning probability of discovery it is clear that we can make the prediction problem arbitrarily hopeless.
We will therefore need extra assumptions. The one which will play a central role for us it the following.
Definition 21.
We say that $\mu$ is of $B$-bounded arity and write $\mu\prec B$ if $\mu(C)=0$ for all $C\in 2^{\mathcal{S}}$ such that $|C|>B$. We say that $\mu$ is of bounded arity if it is of $B$-bounded arity for some $B\in\mathbb{Z}_{\geq 1}$.
In practice $B$ may or may not be known. For example, the genome fragments consist of $B=100$ base pairs. Phone calls are between $B=2$ people. The card game Magic the Gathering (MTG) has packs of $B=15$ cards. On the other hand, when collecting butterflies $B$ is not known, in so far that it may be said to exist at all.
Remark 22.
It is natural to ask whether one could get away with a weaker assumption than bounded arity. Perhaps the most natural would be to consider the mean size of the sampled sets. However, by tuning the probability of discovering the cave and the number of species it contains appropriately one sees that constructions like Example 20 still cause issues and so the mean set size being small is not enough. See Remark 48 for discussion of a weakening in the far-future regime.
Under bounded arity $\delta(t)$ may be controlled similarly to how it was before.
Proposition 23.
If $\mu\prec B$ then,
$\delta(t)\leq r^{2}\mathbb{E}[S_{t}]+\mathbb{E}[S_{t,T}]\leq Br(r+1)t.$
Remark 24.
[…]
Bounded arity is enough to get excellent control of $\epsilon(t)$.
Theorem 25.
If $\mu\prec B$ then,
$\epsilon(t)\leq r(r+1)(B-1)\mathbb{E}[S_{t}].$
Proof sketch.
One imagines constructing $\mu$ by the following procedure. Start by giving zero mass to all sets. Order the sets in some fixed but arbitrary way. At the $n$:th step, slowly add mass to the $n$:th set. Computing the derivatives with respect to the added probability mass one sees that the left hand side grows slower than the right hand side and therefore ends up smaller once the construction is done. ∎
Combining Theorem 19, Proposition 23, Theorem 25, Lemma 39 and Remark 24 one obtains the following.
Corollary 26.
For $r\leq 1$ the GTE is minimax-optimal to within a factor $r+1\leq 2$ also for the generalized unseen species problem with $\mu$ of $B$-bounded arity for any $B$.
However, this corollary undersells Theorem 25 which really establishes that $\epsilon(t)$ has the expected rate not just in $t$ but in $\mathbb{E}[S_{t}]$.
[…]
Proposition 27.
Let $\hat{\epsilon}(t):=\sum_{x,y}(r^{2N_{x\cap y}}-(-r)^{N_{x\cap y}})(-r)^{N_{x\otimes y}}$. Then,
$\epsilon(t)=\mathbb{E}[\hat{\epsilon}(t)].$
Unfortunately some numerics suggest that $\hat{\epsilon}(t)$ is not particularly well-concentrated.
As expressed in the following proposition, if we assume a small multiplicative difference between the unions and the intersections we can take the number of pairs co-observed precisely once as a conservative estimator of $\epsilon(t)$.
Proposition 28.
If $\left(1+\frac{r^{2}}{1+r}\right)M_{x\cap y}<M_{x\cup y}$ for all $x,y\in\mathcal{S}$ then
$\epsilon(t)\leq r(r+1)\mathbb{E}\left[\sum_{x,y}1_{N_{x\cap y}=1}\right].$
Proposition 29.
Let $A_{x,y}$ be the event that $x,y$ are co-discovered, namely the event that the first set which contains either $x$ or $y$ contains both. Then,
$\epsilon(t)\leq r(r+1)\sum_{x,y}\frac{M_{x\cap y}}{M_{x\cup y}}=r(r+1)\mathbb{E}[\sum_{x,y}1_{A_{x,y}}].$
Note that $\sum_{x,y}1_{A_{x,y}}$ is not $\mathcal{F}_{t}$-measurable but that since likely most co-discoveries occur fairly early the number of co-discoveries up to time $t$ can serve as a reasonable approximation. Proposition 29 justifies the following heuristic: If the number of co-discovered pairs is small compared to the number of discovered species, then the dependence is negligible.
4 Intermediate Future
4.1 Classical Setting
4.1.1 Algorithmically Constructing Linear Estimators
Following previous authors we aim to construct linear estimators suited to $r>1$. These will not be a-temporal as to avoid being subject to Theorem 15 via Proposition 10. Instead of trying to find an interpretable $H$ à la [20] we aim to choose one algorithmically. At face value this is not a feasible optimization problem as one would need to optimize over both $H$ and $\mu$. However, we give the following theorem which alleviates the problem. We denote the exponential generating function of $H$ (understood as a sequence) by
$g_{H}(x):=\sum_{i=1}^{\infty}\frac{H_{i}}{i!}x^{i},$
and let $H^{2}$ denote the element-wise square of $H$.
Theorem 30.
For $r,t>0$ and a linear estimator $\hat{S}_{t,T}^{(H)}$ set
$G_{H}:=\left(\sup_{p\in[0,1]}\frac{e^{-pt}}{p}|1-e^{-rpt}-g_{H}(pt)|\right)^{2}$ $+\sup_{q\in[0,1]}\frac{e^{-qt}}{q}(1-e^{-rqt}+g_{H^{2}}(qt)).$
Then,
$\frac{1}{82}G_{H}\leq\sup_{\mu}\mathbb{E}[|S_{t,T}-\hat{S}^{(H)}_{t,T}|^{2}]\leq G_{H}.$
Proof sketch.
Imagine trying to construct a $\mu$ so that the bias is large. In species $s$ the bias is $b_{s}:=\mathbb{E}[1_{N_{s}=0}1_{N_{s}^{\prime}>0}-H(N_{s})]=e^{-p_{s}t}(1-e^{-rp_{s}t}-g_{H}(p_{s}t))$. One can therefore interpret $\frac{e^{-p_{s}t}}{p_{s}}|1-e^{-rp_{s}t}-g_{H}(p_{s}t)|$ as the ”cost effectiveness” of ”spending” $p_{s}$ to cause bias. Spending the total probability mass of $1$, the best cost-effectiveness upper bounds the maximal achievable bias. The variance is bounded in a similar way, giving the upper bound in the theorem. For the lower bound, if $p^{*}$ and $q^{*}$ (the points at which the suprema are attained) happen to be reciprocals of integers then one can either realize the bias bound or the variance bound by taking a uniform distribution on $\frac{1}{p^{*}}$ or $\frac{1}{q^{*}}$, symbols, respectively. Interpolating these gives a lower bound of $\frac{1}{2}G_{H}$. As they may not be reciprocals of integers significant additional care is needed. ∎
We do not believe that the lower bound is sharp. If one is willing to restrict oneself to ”reasonable” $H$, which make a genuine attempt at the prediction problem, one can no doubt improve the lower bound substantially as such $H$ tend to have a uniformly bounded $b_{s}$ and small $q^{*}$ and $p^{*}$ (which are hence well-approximated by reciprocals of integers). The upper bound is sharp in that it is an equality for the GTE.
In light of Theorem 30 it is very natural to seek an $H$ that minimizes $G_{H}$. This turns out to be a well-behaved optimization problem. Critically, $G_{H}$ is convex and even strictly convex away from $q^{*}=0$ in the following sense.
Proposition 31.
Let $\theta\in(0,1)$. For any sequences $H^{(1)},H^{(2)}$,
$G_{\theta H^{(1)}+(1-\theta)H^{(2)}}\leq\theta G_{H^{(1)}}+(1-\theta)G_{H^{(2)}}.$ (3)
Moreover, if $H^{(1)}\neq H^{(2)}$ and $\theta$ are such that the $q^{*}$ associated to $\theta H^{(1)}+(1-\theta)H^{(2)}$ is non-zero and $G_{\theta H^{(1)}+(1-\theta)H^{(2)}}<\infty$, then strict inequality holds in (3).
Convexity means we can hope to find a good $H$ by (sub-)gradient based methods. Moreover, the following theorem guarantees that there is something to converge to.
Theorem 32.
The functional $G_{H}$ admits a minimizer. That is, for all $r,t>0$ there exists a $H^{*}$ such that $G_{H^{*}}\leq G_{H}$ for all other sequences $H$. Moreover there exists an $r_{0}$ such that for all $r>r_{0}$ there exists some $t_{0}(r)$ such that for all $t>t_{0}$ the minimizer is unique.
Proof sketch.
The proof of existence is via the direct method in the calculus of variations. $G_{H}$ is a rather natural functional on the Hilbert space $\mathcal{H}_{t}$ with inner product $\langle H^{(1)},H^{(2)}\rangle_{t}:=\sum_{i=1}^{\infty}H^{(1)}_{i}H^{(2)}_{i}\frac{t^{i}}{i!}$ and so can be shown to be lower semi-continuous. This combines with convexity to give weak lower semicontinuity, the key requirement for the direct method.
For uniqueness one needs an argument that optimizers don’t occur on the part of the space where strict convexity fails. The spirit of this argument is that $q^{*}=0$ implies a bound on $g_{H^{2}}(qt)$, which forces the coefficients to be small enough that a large bias is unavoidable. ∎
Remark 33.
Our uniqueness proof relies on the fact that for $r$ large enough there exists an $x$ such that
$rx-\sqrt{xe^{x}(r+r^{2})-1+e^{-x}-r^{2}x}\sqrt{e^{x}-x-1}>1-e^{-rx}.$
Numerically this seems to be true down to $r_{0}\approx 2.88$. For smaller $r$ different ideas would be needed, if uniqueness holds at all.
From Theorem 30 we obtain the following corollary.
Corollary 34.
Let $\hat{S}_{t,T}^{(H^{*})}$ be the linear estimator associated to a (the) minimizer of $G_{H}$. Then for any $r,t>0$ and any other linear estimator $\hat{S}_{t,T}^{(H)}$ we have
$\sup_{\mu}\mathbb{E}[|S_{t,T}-\hat{S}_{t,T}^{(H^{*})}|^{2}]\leq 82\sup_{\mu}\mathbb{E}[|S_{t,T}-\hat{S}_{t,T}^{(H)}|^{2}].$
Since the GTE and SGTEs are linear, this in particular means that $\hat{S}_{t,T}^{(H^{*})}$ achieves (at least) their rates and hence obtains the optimal rate for $r\leq 1$ and is minimax pseudo-optimal for $r>1$.
4.1.2 Implementation
Of course on a computer we can only minimize over a finite dimensional space and so for the purposes of producing the figures in the present paper we restrict ourselves to $H$ such that $H_{k}=0$ for all $k$ larger than $10$. To minimize $G_{H}$ we approximate the suprema using softmax and use a gradient based method (L-BFGS-B) over a gradually refining grid. Code for the solver as well as the code used to generate all the figures in the paper, is available on our GitHub 333github.com/EdwardEriksson/Unseen-Species. We will sometimes abuse terminology/notation and speak of the finite-dimensional approximation as if it was $\hat{S}_{t,T}^{(H^{*})}$.
4.1.3 Comparing Worst Case Guarantees
Once our solver finds an $H$ we evaluate $G_{H}$ on a very fine grid to get a bound on the mean squared error. In particular, one need not believe that our optimization procedure found the optimal $H$ to trust the bound on the MSE. It suffices that we can evaluate $G_{H}$ accurately for the found $H$. We also evaluate $G_{H}$ for the SGTE, hence obtaining a MSE guarantee for it. We show the worst case guarantees we can give by this method against those for the SGTE with the analysis of [20] in Figure 1.
Figure 1: Comparison of the worst case RMSE of the SGT (as guaranteed by the work of [20]) and our estimator. For the SGT we choose the truncation distribution to be the binomial distribution with the suggested parameter choice of [20]. We also evaluate $G_{H}$ for the SGT, obtaining MSE guarantees for the SGT from our analysis. Note that the top plots use log-log scales.
It is clear that our worst case guarantees are much better and that most of the improvement is not from the analysis being better but rather from our choice of estimator. On the other hand, we find that the suprema are attained by very small $p^{*},q^{*}$, which suggests that distributions with many rare species are binding on the optimization problem. This is natural as we are considering an absolute rather than a relative error and $\mu$ with few rare species are ”easy” in this metric. This means that we should not expect much if any improvement on the SGTE for $\mu$ which have few rare species. We find that this is true empirically. See the uniform distribution and MTG parts of Figure 2 and surrounding discussion. Relatedly, we note that Theorem 30 can easily be adapted to the assumption of lower bounded probabilities which is a common assumption in support size estimation [23]. Specifically, if one restricts the supremum to be over $\mu$ with no species less probable than $p_{0}$ for some $p_{0}>0$, then the upper bound in Theorem 30 holds with the suprema restricted to $p_{0},1]$. Thus one can construct, using essentially the same methodology, estimators which are more well-suited to the regime of few rare species.
4.1.4 Central Limit Theorem for Linear Estimators
For a linear estimator we suggest estimating the variance by
$\hat{V}^{(H)}_{t}:=\hat{S}_{t,T}^{(H)}+\sum_{s\in\mathcal{S}}H(N_{s})^{2}.$
With this variance proxy we give a fairly generic CLT for linear estimators.
Theorem 35.
Let $\hat{S}_{t,T}^{(H)}$ be a linear estimator, not necessarily atemporal, and let $r(t)$ possibly vary in time. Suppose that $\mathbb{E}[\hat{S}^{(H)}_{t,T}]\geq 0$ and $|b_{s}|\leq c$ uniformly in time for some constant $c$. Suppose also that
[…]
Then,
$\frac{S_{t,T}-\hat{S}_{t,T}^{(H)}-\mathbb{E}[S_{t,T}-\hat{S}_{t,T}^{(H)}]}{\sqrt{\hat{V}_{t}}}\rightarrow\mathcal{N}(0,1).$
Remark 36.
Theorem 35 can be seen to imply Theorem 17 by noticing that the GTE has $b_{s}\equiv 0$ and that for $r\leq 1$, $H||_{\infty}\leq 1$.
Remark 37.
The authors of [20] give upper bounds on $b_{s}$ and $H||_{\infty}$ for their estimators, so for suitable $\mu$, $r(t),$ they will satisfy the assumptions of the theorem.
4.2 Generalized Setting
Let $R_{t}=\sum_{x,y}1_{N_{x\cap y}=0}1_{N_{x\cap y}^{\prime}>0}$ be the number of pairs which have not been co-observed but which are about to be and let $Q_{t}=\sum_{x,y}1_{N_{x\cap y}>0}$ be the number of pairs that have been co-observed.
Theorem 38.
Let $\hat{S}_{t,T}^{(H)}$ be a linear estimator with $H$ chosen such that, with $N,N^{\prime}$, independent Poisson random variables with means $\lambda,r\lambda$ there exists a function $f$ such that $\mathbb{E}[H(N)-1_{N=0}1_{N^{\prime}>0}]\leq f(\lambda,r)$ and $f$ is non-decreasing in $\lambda$. Then,
[…]
We also provide the following estimates.
Lemma 39.
If $\mu\prec B$ then,
$\mathbb{E}[S_{t}]$ $\leq Bt,\hskip 10.00002pt$ $\mathbb{E}[S_{t,T}]$ $\leq Brt,\hskip 10.00002pt$ $\mathbb{E}[R_{t}]$ $\leq rB(B-1)t,\hskip 10.00002pt$ $\mathbb{E}[Q_{t}]$ $\leq B(B-1)t.\hskip 10.00002pt$
The combination of Theorem 38 and Lemma 39 means we can bound the error in using linear estimators on incidence data in quite some generality. In particular, since [20] shows that for SGTEs, $H||_{\infty}\leq\mathbb{E}[r^{L}]$ and $\mathbb{E}[1_{N_{s}=0}1_{N_{s}^{\prime}>0}-H(N_{s})]\leq(1-e^{-rM_{s}t})\xi_{L}(r)$ (see their paper for the definition of $\xi$) which is increasing in $M_{s}t$ and $\sum_{s\in\mathcal{S}}(1-e^{-rM_{s}t})\xi_{L}(r)=\mathbb{E}[S_{T}]\xi_{L}(r)$ we have the following corollary.
Corollary 40.
The SGTEs of [20] including the Efron-Thristed estimator, achieve the same rate on the generalized unseen species problem with $\mu\prec B$ for any $B$ as they do in the classical unseen species problem. Since allowing $\mu\prec B>1$ only makes the problem harder this means that the SGTEs are pseudo-optimal in the bounded arity version of the generalized problem with the same rate gap as in the classical problem.
5 Distant Future
5.1 Classical Setting
Recall that in the distant future regime the unseen species problem essentially reduces to the problem of estimating $\alpha$ and that this may be done in different ways. We give a CLT-like result for an unseen species estimator obtained in this way. We write $f(x)\sim g(x)$ to denote $\lim_{x\rightarrow 0}\frac{f(x)}{g(x)}=1$.
Theorem 41.
Let $\nu(x)\sim cx^{-\alpha}$ for some constants $c>0$ and $\alpha\in(0,1)$ and let $r\in\mathbb{R}_{>0}$ be fixed. Let $\hat{\alpha}$ be any estimator of $\alpha$ which converges to $\alpha$ in probability. Let $\hat{S}_{t,T}:=S_{t}((1+r)^{\hat{\alpha}}-1)$ be the induced estimator for $S_{t,T}$. Let $\sigma^{2}:=\Gamma(1-\alpha)[(2^{\alpha}-1)(1+r)^{\alpha}-2(1+r)^{\alpha}((1+\frac{1}{1+r})^{\alpha}-1)+2^{\alpha}-1]$. Then,
$\frac{\hat{S}_{t,T}-S_{t,T}-\mathbb{E}[S_{t}](1+r)^{\hat{\alpha}}+\mathbb{E}[S_{T}]}{t^{\frac{\alpha}{2}}}\rightarrow\mathcal{N}(0,\sigma^{2}).$
Moreover, if $\hat{\alpha}$ estimates $\alpha$ at a better rate than $\frac{1}{\log(t)}$ then the statement continues to hold upon replacing the $\alpha$ in the denominator by $\hat{\alpha}$. Note also that by the continuous mapping theorem $\hat{\sigma}\rightarrow\sigma$ where $\hat{\sigma}$ is the plug-in estimator using $\hat{\alpha}$.
We also give the following straightforward proposition, aimed at giving a quantitative idea of the size of the bias-like term in the above.
Proposition 42.
If $|\nu(x)-cx^{-\alpha}|\leq f(x)$ for some function $f$ for all $x\in(0,1)$ and some constants $c>0,\alpha\in(0,1)$ then, almost surely,
$|\mathbb{E}[S_{t}](1+r)^{\hat{\alpha}}-\mathbb{E}[S_{T}]|$ $\leq c(1+r)\log(1+r)\Gamma(1-\alpha)t^{\alpha}|\hat{\alpha}-\alpha|,$
$+(1+r)t\mathcal{L}[f](t)+T\mathcal{L}[f](T),$
where $\mathcal{L}[f](t)$ denotes the Laplace-transform of $f$.
5.2 Generalized Setting
We generalize $\nu(x)$ in the natural way, namely by setting $\nu(x):=\sum_{s\in\mathcal{S}}1_{M_{s}>x}.$
Theorem 43.
Let $\hat{\alpha}=\frac{\phi_{1}}{S_{t}}$. Fix $\epsilon>0$. If there exist constants $c,\alpha,K$ such that $|\nu(x)-cx^{-\alpha}|\leq Kx^{-\frac{\alpha}{2}}$ for all $x\in(0,1)$ then there exists a constant $C(\epsilon,c,\alpha,K)$ so that
$\limsup_{t\rightarrow\infty}P\left(|\alpha-\hat{\alpha}|>\frac{C}{t^{\frac{\alpha}{2}}}\right)<\epsilon.$
This is a factor $\sqrt{\log(t)}$ improvement on Theorem 2 of [17]. We assumed a slightly better power law approximation, but if we make only the weaker assumption we recover their rate, so our result is strictly better. We note that this seemingly marginal rate improvement makes the difference between the bound on the bias-like term in Proposition 42, after the division by $t^{\frac{\alpha}{2}}$, staying bounded or going to infinity.
Remark 44.
Because our suggested estimator for alpha is invariant under the map $\boldsymbol{\phi}\rightarrow k\boldsymbol{\phi}$ the induced estimator $\hat{S}_{t,T}$ is $\phi$-homogeneous while that of [17] is not.
Let $S_{t}^{(i)}=\sum_{j=i}^{\infty}\phi_{j}$ be the number of species seen at least $i$ times and note that $S_{t}=S_{t}^{(1)}$. We give the following concentration inequality.
Lemma 45.
Let $\mu\prec B$. Then,
$P(S_{t}^{(i)}-E[S_{t}^{(i)}]\geq-z)$ $\leq\exp\left(-\frac{z^{2}}{2((B-1)i+1)E[S_{t}^{(i)}]}\right),$
$P(S_{t}^{(i)}-E[S_{t}^{(i)}]\geq z)$ $\leq\exp\left(-\frac{z^{2}}{2((B-1)i+1)\mathbb{E}[S_{t}^{(i)}]+\frac{2}{3}((B-1)i+1)z}\right).$
Proof sketch.
An application of the work of [3] on concentration via size-biased couplings for configuration models. A bounded size-biased coupling can be constructed since $\mu$ is of bounded arity, which implies concentration. ∎
Remark 46.
Note that Lemma 45 does not require a power law approximation to hold.
Corollary 47.
Lemma 45 implies that for $\phi_{i}=S_{t}^{(i)}-S_{t}^{(i+1)}$ we have,
$P(|\phi_{i}-\mathbb{E}[\phi_{i}]|>z)$ $\leq\exp\left(-\frac{z^{2}}{8((B-1)i+1)E[S_{t}^{(i)}]}\right)$
$+\exp\left(-\frac{z^{2}}{8((B-1)i+1)\mathbb{E}[S_{t}^{(i)}]+\frac{4}{3}((B-1)i+1)z}\right)$
$+\exp\left(-\frac{z^{2}}{8(Bi+B-i)E[S_{t}^{(i+1)}]}\right)$
$+\exp\left(-\frac{z^{2}}{8(Bi+B-i)\mathbb{E}[S_{t}^{(i+1)}]+\frac{4}{3}(Bi+B-i)z}\right).$
Remark 48.
As described on page 3285 of [3], the authors of [12] show that the bounded arity assumption in Lemma 45 can be relaxed to $\mu$ being such that there exists a size-biased coupling of $S_{t}$, which we denote by $S_{t}^{(b)}$, so that there exists a $p\in(0,1]$ such that for all $x\in\mathbb{Z}_{\geq 0}$,
$P(S_{t}^{(b)}-S_{t}\leq B|S_{t}^{(b)}\geq x)$ $\geq p,$
$P(S_{t}^{(b)}-S_{t}\leq B|S_{t}^{(b)}\leq x)$ $\geq p.$
As previously mentioned, $B$-bounded arity allows the construction of a size-biased coupling so that $|S_{t}^{(b)}-S_{t}|\leq B$ almost surely.
Remark 49.
In the case $B=1$, the variance factors in Lemma 45 are never better than Proposition 3.4 of [4] and sometimes worse. For instance, for $i=1$ our result has variance factor $\mathbb{E}[S_{t}]$ while theirs has $\mathbb{E}[\phi_{1}]$.
The following is the main theorem of the section.
Theorem 50.
Let $\hat{S}_{t,T}:=S_{t}((1+r)^{\frac{\phi_{1}}{S_{t}}}-1)$, $p\in(0,1)$ and q:=1pq:=1-p. Suppose that $\mu$ is such that for some constants $c>0$ and $\alpha\in(0,1)$ we have $|\nu(x)-cx^{-\alpha}|\leq f(x)$ for all $x\in(0,1)$. Let
$d(z)$ $:=\frac{pzc\Gamma(1-\alpha)t^{\alpha}(1+r)^{-\alpha}}{(1+2\ln(1+r))c\Gamma(1-\alpha)t^{\alpha}+pz(2-\alpha)\ln(1+r)(1+r)^{-\alpha}}.$
Then for $z<\frac{1}{p}(1+r)^{\alpha}(1+2\log(1+r))c\Gamma(1-\alpha)t^{\alpha}$,
$P(|\hat{S}_{t,T}-S_{t,T}|>z)$ $\leq\exp\left(-\frac{(qz-T\mathcal{L}[f](T))^{2}}{2B\mathbb{E}[S_{T}]}\right)$
$+\exp\left(-\frac{(qz-T\mathcal{L}[f](T))^{2}}{2B\mathbb{E}[S_{T}]+\frac{2}{3}B(qz-T\mathcal{L}[f](T))}\right)$
$+\exp\left(-\frac{(d(z)-t\mathcal{L}[f](t))^{2}}{2B\mathbb{E}[S_{t}]}\right)$
$+\exp\left(-\frac{(d(z)-t\mathcal{L}[f](t))^{2}}{2B\mathbb{E}[S_{t}]+\frac{2}{3}B(d(z)-t\mathcal{L}[f](t))}\right)$
$+\exp\left(-\frac{(d(z)+t^{2}\mathcal{L}[f]^{\prime}(t))^{2}}{(4B-2)\mathbb{E}[S_{t}^{(2)}]}\right)$
$+\exp\left(-\frac{(d(z)+t^{2}\mathcal{L}[f]^{\prime}(t))^{2}}{(4B-2)(\mathbb{E}[S_{t}^{(2)}])+\frac{4B-2}{3}(d(z)+t^{2}\mathcal{L}[f]^{\prime}(t))}\right).$
Proof sketch.
The application of Lemma 45 along with a lemma giving bounds for a certain transcendental equation arising from the sensitivity to deviations in the random variables. ∎
The lemma was envisioned in conversation with Anthropic’s Claude and formalized in LEAN by Harmonic’s Aristotle.
Define the normalized error metric $\ell_{\alpha}(S_{t,T},\hat{S}_{t,T}):=\frac{|S_{t,T}-\hat{S}_{t,T}|^{2}}{(rt)^{2\alpha}}.$
Corollary 51.
If $r(t)$ is such that $\frac{r(t)}{\exp{(\sqrt{t^{\alpha}})}}\rightarrow 0$, $f(x)\leq Kx^{\frac{\alpha}{2}}$ for all $x\in(0,1)$ and $\hat{S}_{t,T}:=S_{t}((1+r)^{\phi_{1}/S_{t}}-1)$ then for every $\epsilon>0$ there exists a constant $C(\epsilon,\alpha,c,K,p)$ such that
$\limsup_{t\rightarrow\infty}P(\ell_{\alpha}(S_{t,T},\hat{S}_{t,T})>\frac{C\log(r)^{2}}{t^{\alpha}})<\epsilon.$
Remark 52.
We thank Alexander Fuchs-Kreiss for adapting the techniques of [17] to give a purely asymptotic argument that the above can in fact be strengthened to $r(t)\lesssim\exp(t^{\frac{\alpha}{2}})$ and also for many valuable conversations regarding the present work. This is a factor $\log(t)$ improvement on Theorem 3 of [17] but under a slightly stronger assumption on the power-law approximation. However, if we make only the weaker assumption we end up with their rate, so our result is strictly better.
6 Experiments
6.1 Task
For $r>1$ we compare the performance of a variety of estimators. Our prescribed task is modeled after the use case in [1]. Following [15], [23] and [20] we consider a person reading Hamlet. They know the total number of words in the play and aim to predict the total number of distinct words from reading an initial portion of the piece. Setting $\hat{S}_{T}:=S_{t}+\hat{S}_{t,T}$ we see that this is an instantiation of the unseen species problem. We study absolute relative error, namely $\frac{|\hat{S}_{T}-S_{T}|}{S_{T}}$. When it makes sense to do so we give both the performance when the data is in the true temporal order and an average over $100$ shuffles to reduce noise. The shuffling is justified in so far that we believe in exchangeability. We prefer this to sampling i.i.d. from the empirical distribution as that underestimates the number of rare species and artificially introduces exchangeability.
6.2 Data
In addition to Hamlet we consider data sets of butterfly collecting [24], messages on a social media network [21] (sub-sampled), genome coverage (SRA code SRX151616) (restricted to the first $10^{5}$ genome locations, sub-sampled), simulated packs of MTG cards [5] and some synthetic data ($N=2000$ samples) from the uniform distribution over either $M=10^{3}$ or $M=10^{4}$ species. For all but the Hamlet data we use the generalized unseen species model. The data is described in more detail in Appendix 2.
6.3 Benchmarks
We compare $\hat{S}_{t,T}^{(H^{*})}$, ($H^{*}$) and our suggested distant future estimator (Ratio-$\alpha$) to several competitors. A naive benchmark is obtained from the null estimator (Trivial). We include the binomial smoothing SGTE with the suggested parameters (SGT) and the power law estimator of [17] (MLE-$\alpha$). We also compare against the method of [14] (Pade-GT[2,3][2,3]) developed for applications in genome coverage. We also implement the support size estimator of [23] (Chebyshev). In their notation we set $k$ to be the corpus length and $c_{0}=0.45,c_{1}=0.5$ as their experimentally chosen values.
Figure 2: Empirical Comparison of Performance of Estimators
6.4 Results
See the result in Figure 2. We see that for the messages data the plots for the true order and the permutation average look totally different. This strongly suggests that exchangeability was severely violated and that the model is therefore not a good fit. One sees that the excellent performance of our estimators on the permutation average does not translate into good predictions under the temporal ordering. This serves as an important reminder that exchangeability is a non-trivial modelling assumption and that applications may require specialized models.
We note that since the unseen species problem resembles support estimation only for large $r$ it is natural to expect the Chebyshev-polynomial-based estimator of [23] to perform poorly for small $r$, possibly explaining the U-shaped error curve. It appears that the Padé-based estimator is not competitive with the other estimators.
We find that our distant future estimator fairly consistently outperforms that of [17] and that in as previously discussed, the worst-case improvements of $\hat{S}_{t,T}^{(H^{*})}$ over the SGT translate to improvements when new discoveries are common, but not when they are rare. Neither the linear nor the power law estimators are consistently better than the others across data sets. No doubt in part due to variations in how good the power law approximation is.